Compute the electric potential from a source charge of 4.31 x

10^3 C if a test charge will be placed 7.11 x 10^−12 m from it.


1
Expert's answer
2021-10-19T15:07:38-0400

The potential can be calculated with the following equation:


"V=\frac1{4\pi\epsilon_0}·\frac{q}{r},\\\space\\
V=\frac1{4\pi(8.85·10^{-12})}·\frac{4.31·10^3}{7.11·10^{-12}}=5.45·10^{24}\text{ V}"